$8000 was invested into 2 accounts yielding 3% and 3.5% interest. The total interest earned on the 2 accounts was $255. How much money was invested at 3%?
Found 2 solutions by ikleyn, addingup:
Answer by ikleyn(52943)

.
    x +      y = 8000,
0.03x + 0.035y = 225.

Solve this system, and you will find the amount x at 3% and the amount y at 3.5%.


Try to do it yourself.

If you still have questions, let me know.
In this case do not forget to refer to the number of this question (# 1012821) in order for I could identify it. 


Good luck!


Answer by addingup(3677)

- Total amount 8,000
- Of that, a certain amount was invested at 3%. We will call that amount x.
- The rest of the money, 8,000-x, was invested at 3.5%
------------------------------------------------------------
0.03x+0.035(8,000-x) = 255
0.03x+280-.035x = 255
-0.005x = -25
x = 5,000 this is how much was invested at 3%